OpenStudy (anonymous):

Solve the following equation in the complex number system. x^(5)+6x^(3)-8x(2)-48=0

OpenStudy (anonymous):

by some miracle this factors as \[(x^2+6) (x^3-8) = 0\]

OpenStudy (jamesj):

So you see by inspection that x = 2 is a root ... and sat73 has given most of this to you now ...

OpenStudy (anonymous):

solution for first factor is straight forward. write \[x^2+6=0,x^2=-6,x=\pm\sqrt{6}i\]

OpenStudy (jamesj):

Now I think you should find the other roots by yrself! ;-)

OpenStudy (anonymous):

for second factor jamesj said one solution is 0, so you can factor as the difference of two cubes to get \[(x-2)(x^2+4x+4)\] and now quadratic formula for second part i will keep quiet

OpenStudy (jamesj):

" jamesj said one solution is 2", yep.

OpenStudy (anonymous):

yes of course i meant one solution is x = 2, not x = 0 !

OpenStudy (anonymous):

typo its 8x^(2)

OpenStudy (anonymous):

both jamesj and i knew you meant \[8x^2\] the solution is right